A college student is employed as a door-to-door newspaper salesman. Historical data suggests that the student sells a subscription to 10% of the homes he visits. Suppose the student begins going door-to-door at dome point in time. Given that 5 subscriptions are sold at the first 22 homes, what is the probability that the 7th subscription is sold at the 32nd home?
Answer #1

Answer:

Given,

p = 10% = 0.10

q = 1 - p

= 1 - 0.10

= 0.90

Consider, prbability 7th subscription sold at 32nd home & 5 subscriptions at 1st 22 homes

i.e.,

5 subscriptions in 10th homes = 10C5*0.10^5*0.90^5

= 252*0.1^5*0.9^5

= 0.0015

Probability that 5 subscriptions in 22 homes = 22C5*0.1^5*0.9^17

= 26334*0.1^5*0.9^17

= 0.0439

Required probability = 0.0015/0.0439

= 0.0342